# Machine Learning Interatomic Potentials for Million-Atom Simulations of Multicomponent Alloys

### 02 Apr 2026 v1

@@ I Introduction
Machine learning interatomic potentials (MLIPs) have transformed atomistic modelling across metals, alloys [8], and complex inorganic materials [22]. Foundational developments began with Behler–Parrinello-type neural network potentials (NNPs) [4] and Gaussian Approximation Potentials (GAP) [3], and were subsequently extended by several influential frameworks, including the Spectral Neighbor Analysis Potential (SNAP) based on bispectrum descriptors, Moment Tensor Potentials (MTPs) as a systematically improvable polynomial formalism [26], Deep Potential for scalable deep-learning-based force fields, and the Atomic Cluster Expansion (ACE) as a systematically organized body-order expansion [12]. More recently, pretrained or universal MLIPs, such as MACE [23], CHGNet [10], eqV2 [19], SevenNet [28], and Orb [25], have further expanded elemental coverage and improved transferability across broad classes of materials. These advances have substantially improved both accuracy and generality. However, many universal models rely on computationally intensive equivariant graph or message-passing architectures, which can make routine long-timescale molecular dynamics (MD) at very large scales considerably more demanding than with lighter-weight MLIP frameworks. As a result, despite their broad applicability and strong predictive performance, current demonstrations of such models have remained focused predominantly on molecular systems or materials simulations at moderate system sizes, typically up to the order of tens of thousands of atoms [38, 39].
@@
The scientific payoff is immediate for high-entropy materials, where chemically frustrated disorder and multiscale microstructures demand very large simulation volumes. High-entropy alloys benefit from supercells that capture short-range order, clustering, dislocation networks, and grain boundaries, often at scales beyond thousands of atoms [44, 9, 7]. High-entropy two-dimensional materials [24] and high-entropy ceramics [43] similarly require both wide elemental coverage and large-scale simulations to probe phase stability, defect thermodynamics, and transport. Related extreme-scale arenas include additive-manufacturing thermal cycles, shock and impact loading, radiation damage cascades, and nanocrystalline grain growth, where million-atom domains are commonly reported or strongly indicated [45, 42].

### 03 Apr 2026 v2

@@ I Introduction
Machine learning interatomic potentials (MLIPs) have transformed atomistic modelling across metals [48, 17], alloys [12], and complex inorganic materials [26, 24, 25, 16]. Foundational developments began with Behler–Parrinello-type neural network potentials (NNPs) [4] and Gaussian approximation potentials (GAP) [3, 47], and were subsequently extended by several influential frameworks, including the spectral neighbor analysis potential (SNAP) based on bispectrum descriptors, moment tensor potentials (MTPs) as a systematically improvable polynomial formalism [30], deep potential (DP) for scalable deep-learning-based force fields [49], and the atomic cluster expansion (ACE) as a systematically organized body-order expansion [11]. More recently, pretrained or universal MLIPs, such as MACE [27], CHGNet [9], eqV2 [21], SevenNet [32], and Orb [29], have further expanded elemental coverage and improved transferability across broad classes of materials. These advances have substantially improved both accuracy and generality. However, many universal models rely on computationally intensive equivariant graph or message-passing architectures, which can make routine long-timescale molecular dynamics (MD) at very large scales considerably more demanding than with lighter-weight MLIP frameworks. As a result, despite their broad applicability and strong predictive performance, current demonstrations of such models have remained focused predominantly on molecular systems or materials simulations at moderate system sizes, typically up to the order of tens of thousands of atoms [42, 43].
@@
The scientific payoff is immediate for high-entropy materials, where chemically frustrated disorder and multiscale microstructures demand very large simulation volumes. High-entropy alloys benefit from supercells that capture short-range order, clustering, dislocation networks, and grain boundaries, often at scales beyond thousands of atoms [51, 8, 12, 7]. High-entropy two-dimensional materials [28] and high-entropy ceramics [50] similarly require both wide elemental coverage and large-scale simulations to probe phase stability, defect thermodynamics, and transport. Related extreme-scale arenas include additive-manufacturing thermal cycles, shock and impact loading, radiation damage cascades, and nanocrystalline grain growth, where million-atom domains are commonly reported or strongly indicated [52, 46].